OpenStudy (anonymous):

Square root of 45x^3y^3

OpenStudy (anonymous):

\[\sqrt{45x ^{3}y ^{3}} = \sqrt{9*5*x*x*x*y*y*y} = 3xy \sqrt{5xy}\]